OpenStudy (anonymous):

Okay, all you math genius' out there, I have this information, but what do I do with it? It has to do with scatter plots and point-slope formula. (Medals.) 5.3=x2 5.2=y2 5.0=y1 5.2=x1 then I put it into the equation y2-y1/x2/x1 and got .2/-.1 now I need to put it into y-y1=m(x-x1)... but the question is, how? I really need help guys!

OpenStudy (anonymous):

Have you found the slope yet? \[m=(y_2-y_1)/(x_2-x_1)\]

OpenStudy (anonymous):

yes, it's \[.2\div-.1\]

OpenStudy (anonymous):

Yes, except it will be positive. x2>x1. This is the same as m = 2

OpenStudy (anonymous):

that is your slope or "m" pick an x,y set and solve ex: (5.3,5.2) and get y-5.2=-2(x-5.3)

OpenStudy (anonymous):

but do I need to solve two separate equations, since i have two sets of numbers?

OpenStudy (anonymous):

Positive 2, yes? And from the problem, I think he's supposed to use (x1, y1) even though it will end up being the same equation.

OpenStudy (anonymous):

No, just a single equation. You have two points, and two points determine a line.

OpenStudy (anonymous):

.2/(-.1)=-2

OpenStudy (anonymous):

Right, but I think that should be .2/+.1

OpenStudy (anonymous):

so I put the x1 and y1 into the equation??? Sorry, I'm just a silly little blonde girl!!! lol... math isn't my strongest suit.... :) I appreciate your help MechE_MST and jabberwock! :)

OpenStudy (anonymous):

Yes, that's right. You're doing fine.

OpenStudy (anonymous):

you got the slope 2 then y-y1 = m(x-x1) y-5 = 2(x-5.2) y-5 = 2x-10.4 y= 2x-5.4 that should be the answer

OpenStudy (anonymous):

Okay Thanks Everyone! :D I FEEL SMARTICALER ALL READY!